# Evaluating Large Language Models in Process Mining: Capabilities, Benchmarks, and Evaluation Strategies

### v2

@@ Abstract
Process Mining Large Language Models (LLMs) Output Evaluation Benchmarking Strategies.
@@ 1 Introduction
The three main contributions of this paper are following. First, building upon prior work [3, 4] proposing textual abstractions of process mining artifacts and an experimental evaluation of LLMs’ responses, the essential capabilities that LLMs must have for PM tasks are derived in Section 3.1. The aforementioned capabilities allow us to narrow down the field of LLMs to those that meet these requirements. Next, evaluation benchmarks for selecting suitable LLMs are introduced in Section 3.2, incorporating both process-mining-specific and general criteria such as reasoning, visual understanding, factuality, and trustworthiness. Finally, we suggest automatic, human, and self-assessment methods for evaluating LLMs’ outputs on specific tasks in Section 3.3, aiming to establish a comprehensive PM benchmark and enhance confidence in LLMs’ usage, addressing potential issues like hallucination.
@@ 3.1 LLMs Capabilities Needed for Process Mining
  • Accepting Visual Prompts: Visualizations in PM, such as the dotted chart and the performance spectrum [15], summarizing process behavior, empower analysts to spot interesting patterns not seen in tables. Interpreting visual prompts is key for semi-automated PM. Large Visual Models (LVMs) use architectures similar to language models trained on annotated image datasets [31]. They perform tasks like object detection and image synthesis, recognizing patterns, textures, shapes, colors, and spatial relations. [^4] GPT-4222https://platform.openai.com/docs/guides/vision and Google Bard/Gemini [^3] https://labelbox.com/product/model/foundry-models/google-gemini-pro-vision/ are popular models supporting both visual and textual prompts
@@
  • Factuality: Hallucination in LLMs refers to instances where the model generates incorrect, misleading, or fabricated information without basis in input data or facts [24]. Factuality is the LLM’s ability to verify its content against established facts or data, preventing fabrication or misrepresentation. Factuality is particularly important for PM tasks such as semantic anomaly detection and root cause analysis and may require using external databases, knowledge bases [19], or internet search [32] to validate responses. For example, in anomaly detection, an identified anomaly (e.g., the activity “Cancel Order” followed by “Deliver Order” in an Order-to-Cash process) can be validated against public information to confirm its legitimacy. Some popular LLMs feature web browsing capabilities, enhancing their factuality by accessing current information. [^5] https://cointelegraph.com/news/chat-gpt-ai-openai-browse-internet-no-longer-limited-info-2021666https://community.openai.com/t/knowledge-base-input-json-vs-csv/494534777https://blog.google/technology/ai/try-bard/
@@ 3.2 Relevant LLMs Benchmarks
  • Visual benchmarks: Interpreting visualizations like dotted charts is vital for PM (c.f. Section 3.1). LLMs must effectively handle queries on such visualizations. Benchmarks like MMBench test models on image-based tasks [17], while MM-Vet evaluates capabilities in recognition, OCR, and other areas [35]. However, these appear not to be entirely suited to analyze process mining visualizations, in which the analysis of the orientation of the lines and the size/color of points is required.
@@
  • Fairness benchmarks: LLMs assess fairness in PM by comparing treatment across groups and detecting biases. DecodingTrust evaluates LLMs’ trustworthiness, including toxicity, bias, robustness, privacy, ethics, and fairness [30].
@@
Direct Provision Code Generation Hypotheses Generation Traditional Domain Knowledge Visual Prompts Text-to-SQL Fairness Hypotheses Generation
Process Description X X X
Process Modeling X X X X X X X
Anomaly Detection X X X X X X
Root Cause Analysis X X X X X X
Ensuring Fairness X X X X X X X
Expl. and Interpreting X X X
@@ 3.3 How to Evaluate LLMs Outputs
  • Human evaluation, pertinent for direct querying and hypothesis generation tasks in LLMs. For direct querying, involving tasks like anomaly detection and root cause analysis, key criteria include recall (the proportion of expected insights identified by the model) and precision (the accuracy of the insights provided). In hypothesis generation, the same criteria of recall and precision are applicable. Moreover, the effectiveness of the feedback cycle, where the output of the model is used to validate the original hypotheses, is a measure for these tasks.
  • Self-evaluation in large language models (LLMs) addresses the hallucination issue, as highlighted by [24]. Various strategies include the chain-of-thought approach, where LLMs elaborate on their reasoning processes, improving explanation quality [33]. Confidence scores allow LLMs to rate their insights’ reliability, encouraging higher quality by discarding less certain outputs [27]. Ensembling, combining results from several LLM sessions through majority voting or confidence evaluation, boosts output accuracy [18]. Self-reflection, where an LLM reviews its output or that of another model, helps catch inaccuracies [25]. In practices like anomaly detection, leveraging confidence scores to filter out uncertain anomalies and ensembling for corroborating detection across sessions enhance reliability.
@@ 4 Conclusion
This paper reviews the current implementations of LLMs in PM and proposes three contributions. First, we identified the required capabilities of LLMs when applied in the context of PM. Then, we described some general-purpose and domain-specific benchmarks proposed in the literature. Eventually, we described different strategies to evaluate the outputs of LLMs on PM tasks. The strategies are important not only to build confidence in using LLMs but also to lay the foundations to build comprehensive PM benchmarks measuring the effectiveness of LLMs on all the implementation paradigms.
As the generative AI field rapidly expands, it is important to note that our discussion focuses on the current model capabilities. We anticipate advancements, such as extracting event logs for PM from process execution videos. However, we expect the requirements outlined in this paper to stay relevant for the foreseeable future. The landscape of benchmarks for PM-on-LLMs tasks is likely to advance, with general-purpose benchmarks encompassing more complex tasks as LLMs improve. Additionally, PM-specific benchmarks should emerge. Nonetheless, the core areas and implementation paradigms are expected to remain stable.

### v3

@@ Abstract
Large Language Models (LLMs) Output Evaluation Benchmarking Strategies.
@@ 1 Introduction
The three main contributions of this paper are summarized in Fig. 1. First, building upon prior work [3, 4] proposing textual abstractions of process mining artifacts and an experimental evaluation of LLMs’ responses, the essential capabilities that LLMs must have for PM tasks are derived in Section 3.1. The aforementioned capabilities allow us to narrow down the field of LLMs to those that meet these requirements. Next, evaluation benchmarks for selecting suitable LLMs are introduced in Section 3.2, incorporating both process-mining-specific and general criteria such as reasoning, visual understanding, factuality, and trustworthiness. Finally, we suggest automatic, human, and self-assessment methods for evaluating LLMs’ outputs on specific tasks in Section 3.3, aiming to establish a comprehensive PM benchmark and enhance confidence in LLMs’ usage, addressing potential issues like hallucination.
@@ 3.1 LLMs Capabilities Needed for Process Mining
  • Accepting Visual Prompts: Visualizations in PM, such as the dotted chart and the performance spectrum [15], summarizing process behavior, empower analysts to spot interesting patterns not seen in tables. Interpreting visual prompts is key for semi-automated PM. Large Visual Models (LVMs) use architectures similar to language models trained on annotated image datasets [31]. They perform tasks like object detection and image synthesis, recognizing patterns, textures, shapes, colors, and spatial relations. [^2] GPT-4 and Google Bard/Gemini are popular models supporting both visual and textual prompts.
@@
  • Factuality: LLM hallucination involves generating incorrect or fabricated information [24]. Factuality measures an LLM’s ability to cross-check its outputs against real facts or data, crucial for PM tasks like anomaly detection and root cause analysis. This may involve leveraging external databases [19], knowledge bases, or internet search [32] for validation. For instance, verifying the sequence Cancel Order” followed by Deliver Order” against public data in anomaly detection. LLMs with web browsing can access up-to-date information, enhancing factuality. [^3] https://cointelegraph.com/news/chat-gpt-ai-openai-browse-internet-no-longer-limited-info-2021
@@ 3.2 Relevant LLMs Benchmarks
  • Visual benchmarks: Understanding PM visualizations, such as dotted charts, is essential (c.f. Section 3.1). LLMs must accurately process queries on these visualizations. MMBench tests models on image tasks [17], and MM-Vet assesses recognition, OCR, among others [35]. Yet, they may not fully meet PM visualization analysis needs, particularly in evaluating line orientations and point size/color.
@@
  • Fairness benchmarks: they evaluate LLM fairness in PM by analyzing group treatment and bias detection. DecodingTrust measures LLM trustworthiness, covering toxicity, bias, robustness, privacy, ethics, and fairness [30].
@@
Direct Provision
Code Generation
Hypotheses Generation
Traditional
Domain Knowledge
Visual Prompts
Text-to-SQL
Fairness
Hypotheses Generation
Process Description
X
X
X
Process Modeling
X
X
X
X
X
X
X
Anomaly Detection
X
X
X
X
X
X
Root Cause Analysis
X
X
X
X
X
X
Ensuring Fairness
X
X
X
X
X
X
X
Expl. and Interpreting
X
X
X
@@ 3.3 How to Evaluate LLMs Outputs
  • Human evaluation is essential for LLM tasks like direct querying and hypothesis generation. For direct querying tasks such as anomaly detection and root cause analysis, important criteria are recall (the model’s ability to identify expected insights) and precision (the correctness of insights). These criteria also apply to hypothesis generation. Additionally, evaluating the feedback cycle’s effectiveness in validating original hypotheses is crucial for these tasks.
  • Self-evaluation in LLMs tackles hallucinations, as noted by [24]. Techniques include chain-of-thought, where LLMs detail their reasoning, enhancing explanations [33]. Confidence scores let LLMs assess their insights’ reliability, discarding uncertain outputs for quality [27]. Ensembling, or using results from multiple LLM sessions, increases accuracy via majority voting or confidence checks [18]. Self-reflection, an LLM reviewing its or another’s output, detects errors [25]. In anomaly detection, using confidence scores to exclude doubtful anomalies and ensembling to confirm detections across sessions improves reliability.
@@ 4 Conclusion
This paper examines LLM applications in PM, offering three main contributions: identification of necessary LLM capabilities for PM, review of benchmarks from literature, and strategies for evaluating LLM outputs in PM tasks. These strategies aim to build confidence in LLM use and establish benchmarks to assess LLM effectiveness across PM implementations.
Our discussion centers on current generative AI capabilities within PM, anticipating advancements like deriving event logs from videos. Despite future enhancements, the criteria discussed here should remain pertinent. Benchmarking for PM tasks on large language models (LLMs) will evolve, including both general and PM-specific benchmarks, yet the foundational aspects and methodologies are expected to stay consistent.